A car is traveling lat 50 km/h on a level road where the coefficent of static friction between its tires and the road is 0.70
Find minimum turning radius of the car
1
Expert's answer
2012-11-19T08:24:56-0500
the minimum radius is the radius when friction force isequal to centrifugal.
Friction force
F_ f = mg*mu,
mu is coefficient of friction.
Centrifugal force is
F_c = m v^2/r
r is minimum radius, So we have
F_f=F_c
mg*mu = m v^2/r
r=v^2/g*mu

r=50^2/(9.8*0.7)= 364.43 m
minimum radius is 364.43 m